

********************************************************************
UARS WEEKLY STATUS REPORT - 12 MAY 2000
********************************************************************
***********
SOLSTICE
***********
We continue to gather enough data to meet all the SOLSTICE science
objectives during this period of reduced tape recorder operations. SOLSTICE
operates normally, collecting solar data on all available orbital tracking
periods and stellar calibration observations whenever possible. Level 2
processing is halted due to bugs in the OA_SSPP_ATT service. CDHF and CSC
personnel are working on the problem.
Today we say goodbye to one of our longtime team members. Dan Gablehouse is
leaving the SOLSTICE program and LASP to take a position at the National
Center for Atmospheric Research here in Boulder. Dan's leadership,
enthusiasm and hard work will be sorely missed but we wish him all the best
with his new opportunities. Thanks Dan.
***********
SUSIM
***********
SUSIM continues to gather a near normal set of daily solar scans
scheduled for periods of predicted telemetry.
The SUSIM operations team has implemented a new version of the
onboard Instrument Operations Program which does not contain code
to perform occultation scans. Occultation scans were discontinued
when the tape recorder was lost. Now with the reduced memory
requirements, we can use the same IOP for both forward and backward
flying.
***********
ACRIM
***********
Nothing received.
***********
PEM
***********
Nothing received.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
Nothing received.
***********
WINDII
***********
Nothing received.
***********
MPG
***********
The MPG reexamined the figures relating solar beta angle to power
available for instrument operations. There were no significant
changes in the indicated power available for instruments.
The list of UARS publications contained on the UARS Science
homepage was updated.
Routine processing of instrument Daily Activity Plans continues with no
problems.
***********
CDHF
***********
STATUS
----------
The CDHF processed current day data through level 3 for all active
instruments. Additionally, CDHF personnel reprocessed: (R)= Requested
(C) = Complete, (F)= Failed (L) = Days remaining to reprocess.
- All requested reprocessing is complete as of this reporting period
The DAAC has been notified that MLS VERSION 4, UARS DAYS 7 - 2103, Subtype
ICE_SATURAT, LEVEL 3AT now resides in the DAAC Transfer Tables. MLS has
requested this data be made available to the User Community.
The CDHF will load March 2000 current day DAAC data in May 2000.
REHOST:
- Level 0 for UARS days 3156 - 3159 were electronically received and
successfully ingested on the Alpha system. These transfers continued to
test the interface between the Alpha and the DCF LZP1 production system.
This data is also being used for analysis, verification, and comparison of
the same data received on the VAX.
***********
FOT
***********
REPORTING PERIOD
This report covers 29 April, 2000 (Orbit 47088, GMT Day 120) through
05 May 2000 (Orbit 47282, GMT Day 126).
SPACECRAFT OPERATION
The observatory is now in Forward Flight (Southern Hemisphere viewing) and is
performing nominally under two-battery operations. The instrument operational
changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
PEM 123/2221 47236 AXIS-1 On
The beta angle ranged between the angles of 12.5 degrees to 42.8 degrees for
this report period. The beta angle is now increasing toward a maximum angle of
+76.0 degrees on 15 May 2000 (DOY 136).
Spacecraft battery 1, 2, and 3 performance monitoring continues. Battery 1
remains OFF the charge relay. Battery 2 and Battery 3 are currently
maintaining greater than 23.8 V end-of-night (EON) load bus voltage. Battery
temperatures are stable with a temperature delta between Battery 2 and Battery
3 of 1.0 to 1.3 degrees C. Additional battery performance data for each
battery on 28 April 2000 and 05 May 2000 are provided in Appendix D.
The clock error ranged from 14.3 to -12.1 msec during this report period.
Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial excursion of
ALERTS or Out-Of-Limits conditions detected by the FOT for this report. A
summary is listed below and the details of ALERT and Out-Of-Limits occurrences
are listed in Appendix B.
% OUT-OF-LIMITS = 03
% ALERTS = 00
The following real-time command sheets were executed for the listed operational
element(s) during this report period (see Appendix A).
% FOT = 17
% PEM = 01
UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.
% FOT = 02
Two (2) Anomaly Reports remained OPEN at the end of this report period (see
Appendix E).
TRANSPONDER FREQUENCIES:
- Transponder A: 00/091 - 2287.496948
- Transponder B: 00/091 - 2287.499113
GROUND SYSTEM OPERATION
Two (2) operational support problems occurred this period. Anomaly Reports are
included in Appendix E.
The data loss calculations supplied by Data Capture Facility (DCF) from mission
start through UTC Day 00/125 show the observatory data loss to be 295 hours, 38
minutes, 10.561 seconds (an increase of 8 minutes 29.352 seconds since the last
report period).
The increase of 8 minutes 29.352 seconds occurred over two (2) days:
- 5 minutes 25.904 seconds on Day 122 due to an incomplete dump segment,
- 3 minutes 3.448 seconds on Day 123 due to a late acquisition and data hit.
These calculations have been recalculated to include 1999/346 through 2000/006
statistics using scheduled times.
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
The FOT performed numerous tests with the simulator to try to narrow down exact
problem with loading the new TMONs. The FOT received help from Dick Quinn and
Dave Lorenz and are very close to getting to the root of the problem. Once the
TMONs are loaded to the simulator they will be thoroughly tested before they
are presented, approved and uplinked.
The FOT continues to develop an online procedure for safehold should it happen
before the new TMONs are in place or during and OBC halt. This procedure has
yet to be finalized and must undergo peer review before it will be ready.
MEETINGS
The FOT attended the UARS monthly meeting. Spacecraft status was given as well
as an update on safehold work.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw slew is scheduled for June 05, 2000 (DOY 157). This maneuver
will be a Forward to Reverse Flight maneuver (Northern Hemisphere viewing).
Attachments: Appendix A, Real-time Commands Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 05 May 2000
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ -------
47181 120/0608 00-179 NBTR NBTR OPS DAY 120
47185 120/1154 00-183 OBC Clk Adj -20.6 to 0.00 msec/day
47193 121/0052 00-185 OBC Clk Adj 0.00 to 20.6 msec/day
47196 121/0611 00-184 NBTR NBTR OPS DAY 121
47207 121/2357 00-186 NBTR NBTR OPS DAY 122
47211 122/0614 00-186 NBTR NBTR OPS DAY 122
47211 122/0614 00-187 OBC Clk Adj 0.00 to 20.6 msec/day
47218 122/1815 00-188 NBTR NBTR OPS DAY 123
47222 122/2316 00-189 OBC Clk Adj fr 0.00 to -20.6 msec/day
47234 123/1817 00-190 NBTR NBTR OPS D124
47236 123/2221 00-191 OBC Clk Adj fr -20.6 to 0.00 msec/day
47236 123/2221 00-192 AXIS 1 Power ON
47240 124/0431 00-193 OBC Clk Adj fr 0.00 to 20.6 msec/day
47250 124/2000 00-194 NBTR NBTR OPS DAY 125
47265 125/2002 00-195 NBTR NBTR OPS DAY 126
47266 125/2142 00-196 OBC Clk Adj fr 20.6 to 0.00 msec/day
47271 126/0623 00-195 NBTR NBTR NOOPS D126
47279 126/1824 00-197 NBTR NBTR NOOPS D127
______________________________________________________________________________
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 05 May 2000
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
--------- ----- ------------ ------
SOSGTEVLTG Y-HI SOLSTICE GT VOLTAGE 47223
HRPTSPBIMON Y-HI HRDI PULSE BUS MON 47245
HRPTSPBIMON Y-HI HRDI PULSE BUS MON 47247
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
--------- ------------ ------
-None this report period
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 05 May 2000
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ---- -----------------
47185 120/11:54 (29 Apr 00) from -20.6 to 0.00 msec/day
47193 121/00:57 (30 Apr 00) from 0.00 to +20.6 msec/day
47211 122/06:25 (01 May 00) from +20.6 to 0.00 msec/day
47222 122/23:18 (01 May 00) from 0.00 to -20.6 msec/day
47236 123/22:21 (02 May 00) from -20.6 to 0.00 msec/day
47240 124/04:37 (03 May 00) from 0.00 to +20.6 msec/day
47266 125/21:46 (04 May 00) from +20.6 to 0.00 msec/day
_______________________________________________________________________________
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 05 May 2000
28 April 2000 - GMT Day 119 Beta = 12.5 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ----------- ---------- ---------- ------ ------- ------
1 -22.4/-44.8 -3.68/-4.94 +0.0/+0.0 15.8 0.000 0.0
2 +33.6/-11.2 +5.47/+3.87 +30.4/-14.4 24.3 1.022 20.5
3 +33.6/-5.6 +4.51/+2.60 +29.6/-15.2 24.2 1.020 20.1
05 May 2000 - GMT Day 126 Beta = 42.8 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ----------- ---------- ----------- ------ ------- ------
1 - 5.6/-33.6 -3.05/-4.62 +0.0/+0.0 15.8 0.000 0.0
2 +22.4/-50.4 +5.79/+4.51 +27.6/-16.0 24.5 1.030 20.9
3 +22.4/-33.6 +4.83/+3.23 +27.6/-16.0 24.5 1.028 20.7
V/T LEVEL CHANGES
ORB# TIME LVL FR-TO
---- ----- ----------
no changes V/T 5
APPENDIX E
UARS ANOMALY REPORTS
Week ending 05 May 2000
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
00-007 FOT 47199 121/1113 14170 Generic Late Acquisition
00-019 FOT 47279 126/1824 N/A TAC 6 Crashed
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
00-007 FOT see list April 00 14170 Generic Late Acq.
00-019 FOT 47279 126/1824 N/A TAC 6 Crashed
ANOMALY INVESTIGATION REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
99-064 FOT 44135 10/8-9/99 N/A HRDI Cal Lamp Errors
99-071 FOT 44398 299/1212 21512 TPF Not Receiv 512
SEVNTFW00/18